ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: not given Sanni, Sikiru A., University of Oyo, Nigeria Lubricating characteristics of a vibrating slider bearing This paper describes the lubricating characteristics of a vibrating slider bearing. Results are presented in form of integrals. The effects of the vibrating slider on the transient load capacity, frictional drag, coefficient of friction and the centre of pressure were investigated and found to be considerate. Estimates of parameters enhancing load capacity and frictionless flow are obtained.